Fermi National Laboratory
 

Accelerator Update for November 6, 2006 - November 8, 2006

 

Monday November 6

The day shift began with the Tevatron (TeV) sector E4 cryo system cooldown continuing, with Operation monitoring the Antiproton Source (Pbar) stack (13.9E10) and the Recycler stash (118.9E10), but with Pbar not stacking due to Accumulator injection work, with Linac, Booster, and the Main Injector (MI) operating normally, with the Collider Detector at Fermilab (CDF), and D-Zero (D0) waiting for a store, and with MiniBooNE, and NuMI taking beam.

Operations reported at 8:48 AM that the TeV was cold.

Operations received permission to power the TeV bus at 9:43 AM.

At 9:59 AM, Linac experts began repairs on the 400MeV Debuncher cavity temperature regulation.

Pbar experts began a two-hour Accumulator study at 10:10 AM.

Operations put the TeV into ramp mode at 11:40 AM.

At 3:36 PM, operators began preparing the TeV for an access. Vacuum technicians investigated the A4 beam valve air pressure problem. They came out of access around an hour later.

Pbar resumed stacking at 5:23 PM.

The Recycler completed stashing antiprotons at 8:15 PM.

Tuesday November 7

The midnight shift began with the TeV experts conducting studies, with Operations monitoring a stack (46.4E10) and stash (145.2E10), with Linac, Booster, and MI operating normally, with CDF and D0 waiting for beam, and with NuMI and MiniBooNE taking beam.

At 1:18 AM, Operations prepared the Recycler to stash, but a problem with injection buckets put the stash on hold. Operations called in an expert.

At 2:29 AM, Operations reported that the Recycler lost 100.0E10 of antiprotons due to a low level RF problem.

At 2:34 AM, Operations called in a Linac expert to work on the 400 MeV Debuncher.

At 3:15 AM, the MI-10 kicker tripped off. It reset okay.

Operations put the TeV into shot setup at 8:38 AM.

Operations established store 5052 at 12:04 PM with an initial luminosity of 100.2E30.

At 12:08 PM, Operations reported that a TeV separator at D17 had sparked and caused a loss of luminosity.

The Recycler completed a stash at 4:02 PM.

The Recycler completed a stash around 7:03 PM.

The Recycler completed a stash around 9:30 PM.

Wednesday November 8

The midnight shift began with Operations monitoring store 5052 (32.39E30) and a stash (151.4E10), with Pbar stacking (11.4E10), with Linac, Booster, and MI operating normally, with CDF and D0 using the colliding beams, and with NuMI and MiniBooNE taking beam.

The Recycler completed a stash around 2:55 AM.

At 5:04 AM, Operations reported that a Linac RF station (KRF3) had gone out of phase. Operators applied some mechanical encouragement to the flow meter and the meter resumed working. Beam resumed at 5:30 AM.

The Recycler completed a stash at 6:42 AM.

Operations terminated store 5052 at 7:30 AM.

Operations put the TeV into shot setup at 7:52 AM.

The Plans for Wednesday

The plans for today are to stack, stash, and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/08/2006    email Fermilab